### v1

# LASLiN: A Learning-Augmented Peer-to-Peer Network

Abstract
We introduce a learning-augmented peer-to-peer (P2P) network design that leverages the predictions of traffic patterns to optimize the network’s topology. While keeping formal guarantees on the standard P2P metrics (routing path length, maximum degree), we optimize the network in a demand-aware manner and minimize the path lengths weighted by the peer-to-peer communication demands. Our protocol is learning-augmented, meaning that each node receives an individual, possibly inaccurate prediction about the future traffic patterns, with the goal of improving the network’s performances. In this paper, we strike a trade-off between significantly improved performances when the predictions are correct (consistency) and polylogarithmic performances when the predictions are arbitrary (robustness). Our protocol builds upon a generalization of skip lists to networks.
We have two main contributions. First, we consider the centralized setting and show that the problem of constructing an optimum static skip list network (SLN) is solvable in polynomial time and can be computed via dynamic programming. This problem is the natural demand-aware extension of the optimal skip list problem; in particular, we use it to generate the predictions of our learning-augmented protocol.
Second, we introduce the Uniform P2P protocol which generalizes skip list networks (SLN) by allowing the nodes to have any real-valued height instead of discrete ones. We show that Uniform achieves state-of-the-art performances: logarithmic routing and maximum degree, both with high probability. We then use Uniform to build a learning-augmented P2P protocol in order to incorporate demand-awareness, leading to our main contribution, LASLiN. We prove that the performances of LASLiN are consistent with those of an optimum static SLN with correct predictions, and are at most a logarithmic factor off the state-of-the-art P2P protocols if the predictions are arbitrary wrong. For the special case of highly sparse demands, we show that LASLiN achieves an exponential improvement in case of correct predictions, and asymptotically matches the performances of state-of-theart, demand-oblivious P2P protocols in case of adversarial predictions.
keywords:
learning-augmented algorithms, peer-to-peer networks, skip list graphs
1 Introduction
Peer-to-peer (P2P) protocols are one of the key technological advancements that allowed decentralized and distributed systems to scale. A vast body of work has been devoted to optimize the efficiency of these protocols [19, 27, 3, 16, 17]. Traditionally, the efficiency of a P2P protocol is estimated with two key metrics which are the routing path length between a pair of nodes and the maximum degree of each node in the network. State-of-the-art P2P protocols achieve both a logarithmic routing path length and a logarithmic maximum degree [29, 13].
Up until recently, most work on P2P protocols focused on demand-oblivious worst-case scenarios, where no prior information about the possible future demand is known. However, a new line of work studies demand-aware networks [28, 5, 12] which can adapt their connectivity properties to better fit the current traffic. Noting that most networking traffics are sparse and partly predictable [4], the idea is to detect the few highly active nodes and move them topologically closer to reduce the overall delays. The model includes a demand matrix which indicates the amount of traffic between any two nodes in the network; A goal of a demand-aware protocol is then to minimize the total weighted path length [28].
This paper presents a P2P network design where each node can adapt its set of neighbours to better fit the future traffic demands. We do not assume that the demand matrix is known by the nodes; Rather, we stick with the distributed nature of P2P protocols and assume that each node has access to a local prediction about the future demands — the prediction may only consist of partial information about the future traffic, an integer number in the case of this work. In our model, a node may obtain a new prediction at any time to use at its own discretion, independently from the other nodes. Crucially, the predictions are untrusted and we evaluate our P2P protocol in the learning-augmented setting with two metrics: consistency and robustness [20]. The consistency of the algorithm is its performance when the predictions are always perfect, its robustness is its performance when the predictions are worst possible.
1.1 Our Setup
P2P Networks
A P2P protocol is an algorithm distributed over multiple nodes that provides routing facilities between its member nodes. On top of a physical, static network, the P2P protocol keeps track of local routing information stored inside each node.
This paper studies P2P network protocols in which each node $u$ holds a height variable $h_{u}$. In addition, we assume that each node holds a unique identifier (id) which defines a total order at any time: for any two different nodes $u$ and $v$, either $u<v$ or $v<u$. A Skip List Network (SLN[21] is a P2P design based on the well-known skip list data structure [25] which establishes the routing informations solely based on the heights and ids of the nodes. In a SLN, two nodes $u$ and $v$ form an edge, i.e., store each other’s id, if and only if there is no node $x$ such that $u<x<v$ and $h_{x}\geq\min(h_{u},h_{v})$ — see Figure 1.
P2P Networks Constraints
Following the overall consensus [13, 23] on P2P networks, we restrict our focus to P2P protocols with the following properties: (1) any node should be able to join or leave the network at any time, (2) the routing should be local, i.e., each new hop through the network is decided solely by one node, and (3) the routing should be oblivious, i.e., the id of the destination node is sufficient to route the message.
Stated the above requirements, we want the network to still work at scale when the number of nodes in the network becomes large. Noting $n$ the number of nodes at the considered time, we aim at networks with:
  • small routing path lengths. Given any two nodes $u$ and $v$, at all time, the routing path length between $u$ and $v$ is ideally logarithmic in $n$ with high probability.
  • small degrees. At all time, the number of known nodes for any node, i.e., the maximum degree, is ideally logarithmic in $n$ with high probability.
All the P2P protocols described of this paper respect those two criteria. To evaluate the performances of our P2P protocols, we compare them to idealized P2P protocols with the restriction that they need to have polylogarithmic path length and maximum degree (see Definition 2.4).
Traffic Demand
At a given time, we assume that the traffic demand between the nodes is modelled with a demand matrix $W\in M_{n}(\mathbb{R})$, where $W(u,v)$ is the frequency of communications from $u$ to $v$ [28].
Learning-Augmented Protocols
This paper introduces and studies learning-augmented P2P protocols. In the general case, a learning-augmented algorithm is an algorithm that takes an additional input, the prediction, to improve its performances. Crucially, the prediction is untrusted. This means that, with regard to the other inputs of the algorithm, the prediction may be perfect (all things being equal, no other prediction results in a better performance), adversarial, or anything in between.
In our design of learning-augmented P2P protocols, we assume that each node can obtain a prediction at any time, independently from one another. The origin of the prediction is not the focus of this paper; it may come from a centralized server handing out traffic predictions to the nodes, or it may be locally computed by each node based on its own traffic history. In this work, the input prediction is an integer interpreted as the height of the node in the SLN.
1.2 Contributions and Technical Novelty
The contributions of this paper are threefold. First, we formulate the Optimum Static Skip List Network (OSSLN) problem and prove it can be solved in polynomial time, with an algorithm based on dynamic programming (Theorem 3.3). This result extends the previous work of [22] to the case with pairwise demands in line with recent demand-aware results [28]. Later on, this algorithm demonstrates that the task of finding optimal predictions for our learning-augmented P2P protocol is also polynomial.
012312345678
Figure 1: A representation of a Skip List Network (SLN) with eight nodes. The nodes have ids ranging from $1$ to $8$. The height of a rectangle equals the height of its corresponding node; Notice that all heights have integer values as this is an SLN. The double-arrowed, horizontal lines are the edges of the SLN.
0112345678
Figure 2: A continuous SLN with real-valued heights in $(0,1)$.
0123412345678
Figure 3: A Continuous SLN obtained by adding node-by-node the heights of the SLN in Figure 1 and the heights of the Continuous SLN in Figure 2. This is a possible outcome of LASLiN where the predicted heights are those of Figure 1 and the uniform, random heights are those of Figure 2.
01212345678
Figure 4: A SLN with all equal heights, effectively turning the SLN into a simple chained list. This SLN has a linear routing path length. This worst-case performance happens when all the nodes get the same predicted height.
01212345678
Figure 5: The addition of continuous, random heights restores the good performances.
As our second contribution, we present a new P2P design called Continuous Skip List Network (Continuous SLN, Definition 4.1), which generalizes Skip List Networks [6] (SLN, Definition 2.1) and skip lists [25]. We show that a SLN with real-valued heights enables a functioning P2P protocol in that we implement the three basic primitives route, join and leave with state-of-the-art, logarithmic message complexity. We note that all known P2P designs that are based on skip lists [16, 14, 3] historically perceive it as chained lists stacked on each other and explicitly use that discrete nature in the protocol implementation, we show that this is not necessary. We then present a Continuous SLN protocol called Uniform that draws its heights uniformly at random in the interval $(0,1)$ and show that Uniform matches the performances of state-of-the-art P2P protocols with both routing path lengths and maximum degrees lower than $O(\log n)$ with high probability.
As our third contribution, we present a learning-augmented P2P protocol, LASLiN (Learning-Augmented Skip List Network), which combines our Uniform protocol with a predicted SLN to strike consistency and robustness guarantees (Definition 2.8). For each node, LASLiN takes as prediction an integer between $1$ and $\lceil\ln n\rceil$ and sets it as its integer heights at initialization time; it then draws a number in $(0,1)$ uniformly at random and adds it to the integer heights to obtain its final heights as a Continuous SLN. We show that LASLiN is comparable to an optimum static SLN when the predictions are correct ($O(1)$-consistent) and at worst a $\ln n$ factor off a demand-oblivious P2P protocol with arbitrary predictions ($O(\log^{2}n)$-robust). Finally we show that, in the special case of a very sparse demand matrix (see Definition 6.5), the version of LASLiN with binary predictions always outperforms any demand-oblivious P2P protocol with high probability: LASLiN obtains an exponentially better cost with correct predictions while it asymptotically matches the cost of the state-of-the-art P2P protocols for arbitrary predictions (see Theorem 6.6). We provide simulations to assess the performance of LASLiN on Zipfian demand matrices.
2 Model and Metrics
In this section, we formally define our model and the metrics we use to evaluate our P2P protocols.
Definition 2.1 (SLN).
A Skip List Network (SLN) $\mathcal{N}$ is a triple $(V,h,E)$ where:
  • $V$ is a totally ordered set of node ids;
  • $h$ is a height function $h:V\longrightarrow\mathbb{N}^{+}$;
  • $E=\{\{u,v\}\;|\;\forall x\in V\;s.t.\;u<x<v,\;h_{x}<\min(h_{u},h_{v})\}$.
Definition 2.2 (Neighbor and Degree).
We say that two nodes $u$ and $v$ are neighbors in a SLN $\mathcal{N}=(V,h,E)$ if $\{u,v\}\in E$. The degree of a node $u$ is its number of neighbors.
We redefine the greedy, local routing path proper to SLN.
Definition 2.3 (Routing Path).
Let $\mathcal{N}=(V,h,E)$ be a SLN and $u,v\in V$ be two different nodes. The routing path from $u$ to $v$ in $\mathcal{N}$ is the sequence of edges of $E$ $s=e_{1},e_{2}\dots e_{k}$ such that: (1) $u\in e_{1}$, (2) $v\in e_{k}$, and (3) $\forall i\in[k]$, if $u<v$ then $e_{i}=\{x,y\}$ where $y$ is the largest node neighbor of $x$ such that $x<y\leq v$, otherwise if $v<u$ then $e_{i}=\{x,y\}$ where $y$ is the smallest node neighbor of $x$ such that $v\leq y<x$.
More intuitively, the routing path goes to the next neighbor that is the closest to the destination (in terms of ids) without “flying over” the destination.
Definition 2.4 (Ideal P2P Protocol).
We say that a (potentially randomized) P2P protocol is Ideal if, at all times, the routing path length between two nodes and the maximum degree are polylogarithmic on $n$ with high probability, where $n$ is the number of nodes at the considered time. A quantity is polylogarithmic when there exists positive constants $c$ and $d$ such that the quantity is smaller than $d\cdot\log^{c}n$.
Definition 2.5 (Cost).
Let $\mathcal{N}=(V,h,E)$ be an SLN and $W$ be a demand matrix. The cost function is:
$\text{Cost}(\mathcal{N},W)=\sum_{u,v\in V}W(u,v)\cdot d_{\mathcal{N}}(u,v),$
where $d_{\mathcal{N}}(u,v)$ is the routing path length between $u$ and $v$ in $\mathcal{N}$.
Definition 2.6 (Optimum Static Skip List Network (OSSLN) problem).
Given a set of node ids $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\longrightarrow\mathbb{N}^{+}$ that minimizes the cost of the SLN $\mathcal{N}=(V,h,E)$.
Definition 2.7 (Learning-augmented P2P protocol).
A learning-augmented P2P protocol ALG is a P2P protocol whose initialization routine takes an additional parameter $p$, the prediction.
In our model, a node that wishes to update its prediction as it already joined the network simply leaves, re-intializes and joins again.
Definition 2.8 (Consistency and Robustness).
We say that a learning-augmented SLN protocol is
  • $\alpha$-consistent if, for all demand matrices, there exists an assignment of predictions to the nodes such that the expected cost of the protocol is at most $\alpha$ times the cost of an ideal optimum static SLN.
  • $\beta$-robust if, for all demand matrices and for all assignment of predictions, the expected cost of the protocol is at most $\beta$ times the cost of an ideal optimum static SLN.
3 Optimum Static Skip List Network (OSSLN)
Given a set of node ids $V$ and a demand matrix $W$, the objective of the Optimum Static Skip List Network (OSSLN) problem is to find a height assignment $h:V\longrightarrow\mathbb{N}^{+}$ that minimizes the total communication cost (as defined in (2.5)) of the SLN $\mathcal{N}=(V,h,E)$ (as defined in (2.1)):
$OSSLN(V,W)=\min_{h}\;Cost(\mathcal{N},W).$ (1)
In [28], the authors presented an algorithm based on dynamic programming to find the Optimum static distributed Binary Search Tree (OBST) in polynomial time ($O(n^{3})$). This algorithm leverages a crucial property: once a vertex $v$ is chosen as the root of a tree induced by the vertices in a given interval, the subproblems induced by the vertex id intervals to the left and to the right of $v$ can be solved independently. The OSSLN problem exhibits a similar structure, but with crucial difference.
If we define a pivot vertex $x\in[l,r]$, $1\leq l\leq r\leq n,$ such that $h_{x}=\max_{w\in[l,r]}h_{w}$, $x$ also partitions the OSSLN problem into two subproblems (on the intervals $[l,v)$ and $(v,r]$). However, a critical distinction arises when $h_{x}<\max_{w\in[l,r]}h_{w}$, in which case $x$ will not be a part of all the paths between vertex pairs $(u,v)$ $u\in[l,x),v\in(x,r]$. Specifically, for a vertex $x\in[l+1,r-1]$ such that $h_{l},h_{r}>h_{x}$, any path between vertices $(u,v)$ where $u\leq l$ and $v\geq r$ does not include $x$. This changes the optimal substructure of the problem, requiring a more sophisticated dynamic programming approach.
Given a SLN $\mathcal{N}$, we define a path $P_{h}(l,r)$ to be the set of vertices in the routing path between the vertices $l$ and $r$ (as defined in (2.3)), excluding $l$, i.e., $l\notin P_{h}(l,r)$. This is a crucial simplification that allows the problem to be decomposed. Under this assumption, we have $d(l,r)=|P_{h}(l,r)|$. W.l.o.g., we assume that $P_{h}(l,r)=P_{h}(r,l)$.
Next, we define $L(x,h)=\min\{i\leq x\;|\;h_{j}\leq h_{x},\;\forall j\in[i,x]\}$, i.e., the vertex of minimum id between the set of all (consecutive) vertices of height lower or equal to $h_{x}$, to the left of $x$, and, analogously, $R(x,h)=\max\{i\geq x\;|\;h_{j}\leq h_{x},\;\forall j\in[x,i]\}$, i.e., the vertex with maximum id between the set of all vertices of height lower or equal to $h_{x}$, to the right of $x$. A vertex $x\in V$ belongs to a path $P_{h}(u,v)$ if and only if there does not exist any pair of vertices $(u^{\prime},v^{\prime})$ such that $u\leq u^{\prime}<x$, $x<v^{\prime}\leq v$, and (both) $h_{u^{\prime}},h_{v^{\prime}}>h_{x}$.
Finally, given an arbitrary $x\in V$, we define $P_{x}(h)$ as the set of all pairs $(u,v)\in V\times V$, such that $x\in P_{h}(u,v)$.
Lemma 3.1.
If $h$ is a solution for the $OSSLN(V,W)$, then $\forall x\in V$ we have that:
$P_{x}(h)$ $=$ $\{(u,v)\mid u<x\;\wedge\;v\in[x,R(x,h)]\}\;\cup$
$\{(u,v)\mid u\in[L(x,h),x)]\;\wedge\;v>R(x,h)\}$
Proof 3.2.
If $u<x$ and $x\leq v\leq R(x,h)$, then, by definition of $R(x,h)$, $h_{x}\geq h_{v}\geq\min(h_{u},h_{v})$. Since $x$ lies between $u$ and $v$ and there are no edges “above" $h_{x}$ in the interval $[u,v]$ (i.e., $\nexists(w,z)\in E$, s.t. $h_{w}>h_{x}$ and $h_{z}>h_{x}$, $w,z]\subseteq[u,v]$), $x$ must be a part of $P_{h}(u,v)$.
Similarly, if $L(x,h)\leq u<x$ and $v>R(x,h)$, then, again by definition of $L(x,h)$, $h_{x}\geq h_{u}\geq\min(h_{u},h_{v})$. Since $x$ lies between $u$ and $v$ and there are no edges "above" $h_{x}$ in the interval $[u,v]$, $x$ must be a part of $P_{h}(u,v)$
It remains to show that no other pair $(u,v)$ belongs to $P_{x}(h)$. If $u\geq x$, then by definition of $P_{h}(u,v)$, $x\notin P(u,v)$. Otherwise, suppose $u<L(x,h)$ and $v>R(x,h)$. Let $u^{\prime}=L(x,h)-1$ and $v^{\prime}=R(x,h)-1$, by definitions of $L(x,h)$ and $R(x,h)$, we have that $h_{x}<h_{u^{\prime}}$ and $h_{x}<h_{v^{\prime}}$. Since $u\leq u^{\prime}<x<v^{\prime}\leq v$, we have that:
$|P_{h}(u,u^{\prime})\cup\{v^{\prime}\}\cup P_{h}(v^{\prime},v)|<|P_{h}(u,x)\cup\{x\}\cup P_{h}(x,v)|.$
Because $P_{h}(u,v)$ is the shortest path from $u$ to $v$, it follows that $x\notin P_{h}(u,v)$.
Since $d(u,v)=|P_{h}(u,v)|$, we can now rewrite the total cost of an SLN $\mathcal{N}=(V,h,E)$ as:
$\text{Cost}(\mathcal{N},W)$ $=$ $\sum_{u,v\in V}W(u,v)\cdot d(u,v)$
$=\sum_{u,v\in V}\sum_{x\in P_{h}(u,v)}W(u,v)$ $=$ $\sum_{x\in V}\sum_{(u,v)\in P_{x}(h)}W(u,v).$
This means that the total cost is equal to the sum of contributions of all vertices $x\in V$, where the contribution of a vertex $x$ to the total cost is defined as:
$\sum_{(u,v)\in P_{x}(h)}W(u,v)$ (3)
OSSLN with Bounded Height: In the OSSLN with bounded height, the height assignment is of the form $h:V\longrightarrow[1,h_{\max}]$, where $h_{\max}$ is a parameter, typically bounded by $O(\log{n})$ or $O(1)$.
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$ and a height bound $h\in[1,h_{\max}]$. Let $nl$, $1\leq nl\leq l$, denote the leftmost vertex id, such that $\forall v\in[nl,l],h_{v}\leq h$, and, similarly, let $nr$, $r\leq nr\leq n$, denote the rightmost vertex id, such that $\forall v\in[r,nr],\,h_{v}\leq h$. Finally, let $\text{OSSLN}(l,r,nl,nr,h,W)$ denote the minimum total cost contributed by all vertices $v\in[l,r]$.
Theorem 3.3.
If $x\in[l,r]$ is a candidate pivot vertex with height $h\in[1,h_{\max}]$, then the following recurrence relation holds $\forall(nl,l,x,r,nr)$, $1\leq nl\leq l\leq x\leq r\leq nr\leq n$:
$\text{OSSLN}(l,r,nl,nr,h,W)=\min\Bigg{\{}\text{OSSLN}(l,r,l,r,h-1,W),$
$\min_{x\in[l,r]}\Bigg{[}\text{OSSLN}(l,x-1,nl,nr,h,W)+$
$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\Bigg{]}\Bigg{\}}.$
Proof 3.4.
The proof is by strong induction on the size of the interval $l,r]\subseteq[1,n]$:
Note that, by definition, $nl=L(x,h)$ and $nr=R(x,h)$ for any possible $h$. Furthermore, by construction, for an interval induced by $l$, $r$, nlnl, nrnr with bounded height $h$, assigning a pivot vertex with height equal to h1h-1, we have $nl^{\prime}=l$ and $nr^{\prime}=r$.
Base case $(l=r)$: In this case, we have that $(x=l)$ and the set of paths $x$ belongs to is given by Lemma 3.1. By setting $nl=L(x,h)$ and $nr=R(x,h)$ in the proof, the vertex contribution to the total cost is given by:
$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)$ (4)
Induction step: For $(h=1)$, we can disregard the first half of the recursion, where we assign the pivot vertex a height lower than $h$. The total contribution of all vertices to the cost becomes:
$\sum_{x\in[l,r]}\left(\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\right),$
which, from the base case, is precisely what the second half of the recurrence equation will return for any choice of the pivot vertex.
Let $1\leq l<r\leq n$ and $h\in[1,h_{\max}]$. Assume that the recurrence equation is true for all $h^{\prime},l^{\prime},r^{\prime}$ such that $(h^{\prime}<h)\leavevmode\nobreak\ \wedge\leavevmode\nobreak\ (r^{\prime}-l^{\prime})<(r-l)$. From previous step also assume that the equation is true for $(r^{\prime}-l^{\prime})=(r-l)$ with height $(h^{\prime}<h)$.
Since, when assigning the pivot vertex with height lower than $h$, by construction, we have that $(nl=l)$ and $(nr=r$). The minimum cost for the solution for any $nl$ and $nr$ is the minimum between keeping the height of the pivot, $(h_{x}=h)$, or decreasing it, $h_{x}\leq(h-1)$, which is given by the minimum between:
$\min_{x\in[l,r]}\Bigg{[}\text{OSSLN}(l,x-1,nl,nr,h,W)+$
$\text{OSSLN}(x+1,r,nl,nr,h,W)+$
$\sum_{u<x}\sum_{v\in[x,nr]}W(u,v)+\sum_{u\in[nl,x)}\sum_{v>nr}W(u,v)\Bigg{]}$
and OSSLN(l,r,l,r,h1,W)OSSLN(l,r,l,r,h-1,W), which is precisely what we want to prove.
If we pre-compute an auxiliary matrix of aggregate weights ($W^{\prime}(x,y)=\sum_{z\leq y}W(x,z),\forall(x,y),1\leq x\leq y\leq n$) in time $O(n^{3})$, then the above dynamic programming algorithm has complexity $\mathcal{O}(n^{5}\cdot h_{\max})$.
OSSLN with Unbounded height (OSSLNU): We now present a dynamic programming algorithm for the OSSLN problem with unbounded heights, i.e., when the height assignment is of the form $h:V\longrightarrow[1,n]$. We show that, in this particular case, the algorithm’s complexity can be reduced from $\mathcal{O}(n^{5}\cdot h_{\max})$ to $\mathcal{O}(n^{3})$.
A crucial observation is that, for any solution to the OSSLNU problem, there exists a solution of cost less than or equal to that of the former, in which, $\forall[l,r],1\leq l\leq r\leq n$, there is a unique pivot vertex, i.e., a unique vertex whose height is the maximum height in that interval.
Lemma 3.5.
Let $h$ be an optimum solution to $OSSLNU(V,W)$. Then, there must exist a solution $h^{\prime}$ such that $Cost((V,h^{\prime},E^{\prime}),W)\leq Cost((V,h,E),W)$, where, for any interval $l,r],1\leq l\leq r\leq n,$ there is exactly one vertex $x\in[l,r]$ with height $h^{\prime}_{x}=\max_{w\in[l,r]}h^{\prime}_{w}$.
Proof 3.6.
Suppose, for contradiction, that any possible optimal solution contains at least one interval $[l,r]$ with at least two vertices $x,y\in[l,r]$ with $h_{x}=h_{y}=\max(h_{l,r})$ and $x\neq y$.
Define $h^{\prime}$ as follows for any vertex $v\in V$:
$h^{\prime}_{v}=\begin{cases}h_{v}+1&\text{if }v=x\text{ or }h_{v}>h_{x}\\ h_{v}&\text{otherwise}\end{cases}$
Note that, while $h^{\prime}_{x}>h^{\prime}_{y}$, this height assignment is order preserving. If $h_{a}>h_{b}$ then $h^{\prime}_{a}>h^{\prime}_{b}$.
If $x\notin P_{h}(u,v)$, since we preserved the order between $x$ and all vertices taller than $x$, $x$ cannot appear as an obstacle between any such vertices. Analogously, since we preserved the order between all vertices with at most the same height as $x$ there can’t be any new edges between vertices with at most the same height as $x$, so the path is unmodified.
If $x\in P_{h}(u,v)$, we know that:
$P_{h}(u,v)=P_{h}(u,x)\;\cup\{x\}\;\cup P_{h}(x,v).$
From our definition of path, we know that $x\notin P_{h}(x,v)$, and therefore $P_{h}(x,v)=P_{h^{\prime}}(x,v)$. Now, let’s consider a vertex $z$, such that $u\leq z<x$ and $h_{z}>h_{x}$. If such a vertex $z$ does not exist, then we can infer that $P_{h}=P_{h^{\prime}}$. Furthermore, if there does exist such a vertex $z$, then:
$|P_{h^{\prime}}|=|P_{h}(u,z)\;\cup\{x\}\;\cup P_{h}(x,v)|\leq|P_{h}|.$
Since $\text{Cost}((V,h,E),W)=\sum_{(u,v)\in V}W(u,v)\cdot d(u,v)$ and $d(u,v)=|P_{h}(u,v)|$, and any path in $h^{\prime}$ is of at most the same size as in $h$, then, $\text{Cost}((V,h^{\prime},E^{\prime}),W)\leq\text{Cost}((V,h,E),W)$. Hence, the lemma holds.
In OSSLNU, because we are only interested in the relative order of vertex heights, rather than in their absolute values, we have that $(nl=l)$ and $(nr=r)$, $\forall[l,r],1\leq l\leq r\leq n$, and the height parameter $h$ becomes unnecessary. This greatly simplifies the recurrence relation, reducing the complexity to $\mathcal{O}(n^{3})$.
Theorem 3.7.
Consider an interval $[l,r]$, $1\leq l\leq r\leq n$. Let $\text{OSSLNU}(l,r,W)$ denote the minimum total cost contributed by all vertices $v\in[l,r]$. Let $x\in[l,r]$ be a candidate pivot vertex. Define $l$ as the leftmost index such that $\forall x,\in[l,r],\forall v\in[l,x],\,h_{v}\leq h_{x}$, and similarly define $r$ as the rightmost index such that $\forall v\in[x,r],\,h_{v}\leq h_{x}$. Then:
$\text{OSSNU}(l,r,W)=\min_{x\in[l,r]}\Big{(}\text{OSSNU}(l,x-1,W)+$
$\text{OSSNU}(x+1,r,W)+$
$\sum_{u<x}\sum_{v\in[x,r]}W(u,v)+\sum_{u\in[l,x)}\sum_{v>r}W(u,v)\Big{)}.$ (5)
Proof 3.8.
The proof is by induction on the size of the interval $[l,r]$.
Note that, by Lemma 3.5, once a pivot vertex $x\in[l,r]$ is chosen as the highest vertex within the interval $[l,r]$, we can assume, without loss of generality, that all vertices in the optimal solutions to the left subinterval [l,x1][l,x-1] and to the right subinterval $[x+1,r]$ have heights strictly lower than $h_{x}$.
Therefore, for a solution $h$ to $\text{OSSLNU}(l,r,W)$, it follows that $l=L(x,h)$ and $r=R(x,h)$.
Base Case $(l=r)$: In this case, the interval contains a single vertex $x=l=r$. In this case, the cost reduces to the contribution of $x$ to the solution cost, which, from Lemma 3.1, and $L(x,h)=l$ and $R(x,h)=r$ as we demonstrated above, it follows that this expression corresponds exactly to:
$\sum_{u<x}\sum_{v\in[x,r]}W(u,v)+\sum_{u\in[l,x)}\sum_{v>r}W(u,v)$ (6)
Inductive Step: Suppose the recurrence holds for all subintervals of $[l,r]$ of size less than (rl+1)(r-l+1). For an interval $[l,r]$ of size greater than one, we consider a candidate pivot vertex $x\in[l,r]$.
The total cost of placing $x$ as pivot of the interval $[l,r]$ equals to the sum of:
  • The optimal cost for the left subinterval [l,x1][l,x-1],
  • The optimal cost for the right subinterval $[x+1,r]$,
  • The cost of all paths that include $x$, which by Lemma 3.1 is given by the Eq.(6).
Since we minimize this total over all pivot choices $x\in[l,r]$, the recurrence holds.
4 Continuous SLN
This section introduces Continuous SLN, a generalization of SLN.
Definition 4.1 (Continuous SLN).
A Continuous Skip List Network (Continuous SLN) $\mathcal{N}=(V,h,E)$ is an SLN (Definition 2.1) except the height function $h$ is defined over the real numbers, i.e., $h:V\longrightarrow\mathbb{R}$.
We prove that continuous skip graphs support a functioning P2P network by implementing the three primitives: route, join and leave.
Internal Variables.
A node $u$ holds the variables:
  • $u$: the unique id of the node
  • $h_{u}$: the height in $\mathbb{R}$ of node $u$
  • $N_{u}$: the set of all neighbors of $u$
  • $D_{u}$: dictionary mapping a neighbor $w\in N_{u}$ to its height
  • $\text{anchor}_{u}$: the node with an id closest to $u$, either the successor or predecessor of $u$.
Route.
The route algorithm (Algorithm 1) takes the destination node id $v$ and hops from node to node until it finds the destination node, returning $(v,True)$ or reaches a dead end in some node $w\neq v$, return $(w,False)$.
1 // code running at node $u$
2 if $u$ = $v$ then
3   return ($u$, True)
4   else if $u<v$ then
5     $S\leftarrow\{x\in N_{u}|u<x\leq v\}$
6     if $S=\emptyset$ then
7       return ($u$, False)
8       else
9         $w\leftarrow\max S$
10         return $w.\text{route}(v)$
11
12         else
13           $S\leftarrow\{x\in N_{u}|v\leq x<u\}$
14           if $S=\emptyset$ then
15             return ($u$, False)
16             else
17               $w\leftarrow\min S$
18               return $w.\text{route}(v)$
19
Algorithm 1 route($v$)
Join.
The join algorithm (Algorithm 2) inserts the node inside the Continuous SLN based on its anchor local variable. The potential neighbor changes of the other nodes propagate through the network with the find_neighbors algorithm (Algorithm 3).
1
2 // code running at node $u$
3 if $\text{anchor}<u$ then
4   $S\leftarrow\{x\in N_{\text{anchor}_{u}}|\text{anchor}<x\}$
5
6   else
7     $S\leftarrow\{x\in N_{\text{anchor}_{u}}|x<\text{anchor}\}$
8
9     if $S=\emptyset$ then
10       $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
11       Update $D_{u}$
12       else
13         if $\text{anchor}<u$ then
14           $\text{succ}\leftarrow\min S$
15
16           else
17             $\text{succ}\leftarrow\max S$
18
19             $N_{u}\leftarrow N_{u}\cup\text{anchor}_{u}.\text{find\_neighbors}(u,h_{v},-\infty)$
20             $N_{u}\leftarrow N_{u}\cup\text{succ}.\text{find\_neighbors}(u,h_{v},-\infty)$
21             Update $D_{u}$
22
Algorithm 2 join()
1
2 // code running at node $u$
3 if current_height = -$\infty$ then
4   $\text{anchor}_{u}\leftarrow v$
5
6  $N\leftarrow\emptyset$
7   if current_height $h_{u}$ then
8     $N\leftarrow\{u\}$
9     $N_{u}\leftarrow N_{u}\cup\{v\}$
10     Update $D_{u}$
11     if $u<v$ then
12       $N\leftarrow\{x\in N_{u}|x\leq v\text{ or }(h_{v}<D_{u}[x]\text{ and }h_{v}<h_{u})\}$
13       for $x\in N_{u}$ do
14         if $x<u\text{ and }h_{u}\leq D_{u}[x]\leq h_{v}$ then
15           $N\leftarrow N\cup x.\text{find\_neighbors}(v,h_{v},h_{u})$
16
17
18
19          else
20             $N\leftarrow\{x\in N_{u}|v\leq x\text{ or }(h_{v}<D_{u}[x]\text{ and }h_{v}<h_{u})\}$
21             for $x\in N_{u}$ do
22               if $u<x\text{ and }h_{u}\leq D_{u}[x]\leq h_{v}$ then
23                 $N\leftarrow N\cup x.\text{find\_neighbors}(v,h_{v},h_{u})$
24
25
26                 return NN
Algorithm 3 find_neighbors($v$, $h_{v}$, current_height)
Leave.
The leave algorithm (Algorithm 4, in this paper’s appendix) shows how an already-inserted node can leave the network. The potential neighbor changes propagate through the network with the delete_neighbors algorithm (Algorithm 5, in this paper’s appendix).
5 The Uniform SLN Protocol
After discussing the case of static SLN in the previous sections, we now model randomized P2P protocols. This section introduces a randomized, Continuous SLN called Uniform where each node chooses its height by picking a random real number in $(0,1)$ uniformly at random.
Definition 5.1 (Randomized Continuous SLN).
A randomized Continuous SLN is a probabilistic distribution over Continuous SLN.
Protocol 1 (Uniform).
Uniform is a randomized, continuous SLN that sets its heights to a real number in $(0,1)$ chosen uniformly at random.
The rest of this section presents the properties of Uniform with respect to routing path lengths and maximum degrees. We show that a Continuous SLN constructed with Uniform has an expected constant degree and an expected routing path length of $O(\log n)$. Also, we can rule out rare bad outcomes in practice: with high probability, the routing path length is $O(\log n)$ and the maximum degree is $O(\log n)$. Those results directly imply that the three primitive described in the previous section have a logarithmic message complexity with high probability. Interestingly, those results are quite general and hold for a vast family of continuous probability laws, namely, as soon as the probability that two heights are equal is zero.
Theorem 5.2 (Expected Path Length).
The routing path between any two nodes has length smaller than $2\ln n$ on expectation.
Proof 5.3.
Let $u$ and $v$ be two nodes. We note $X$ the random variable equal to the routing path length between $u$ and $v$. We note $X_{\text{up}}$ the number of height increases along the path, likewise $X_{\text{down}}$ is the number of path decreases. It holds that $X=X_{\text{up}}+X_{\text{down}}$. We will give an upper bound of $X_{\text{up}}$, the same upper bound will hold for $X_{\text{down}}$ by symmetry. Let $w$ be a node between $u$ and $v$ in the SLN, we note $1_{w}$ the indicator random variable equal to $1$ if (1) $w$ is part of the path between $u$ and $v$, and (2) if the height increases when hoping to $w$. It holds $X_{\text{up}}=\sum_{w\in[u+1,v]}1_{w}$.
We now compute $P[1_{w}=1]$ for any $\in[u+1,v]. ww$ is part of the increasing path if and only if the height of $w$ is strictly greater than every other heights in [u,w1][u,w-1]. The probability that two uniform random variables are equal is zero, hence there exists exactly one node with maximum height in $[u,w]$. The heights are independent and uniformly distributed, hence $w$ is the node with maximum height with probability 1/(|wu|+1)1/(|w-u|+1):
$\forall w\in[u+1,v],\quad P[1_{w}=1]=\frac{1}{|w-u|+1}$
Noticing that $\mathbb{E}[1_{w}]=P[1_{w}=1]$ and gathering the two previous equalities gives:
$\mathbb{E}[X_{\text{up}}]$ $=\sum_{w\in[u+1,v]}\frac{1}{|w-u|+1}=\sum_{i=2}^{|v-u|+1}\frac{1}{i}\leq\ln n$
The last inequality holds as the harmonic sum until $n$, $1+1/2+\dots+1/n$, is lower than $1+\ln n$. The claimed upper bound on the expected path length follows as the same reasoning can be done for $X_{\text{down}}$.
Theorem 5.4 (Routing Path Length with High Probability).
The routing path between any two nodes has length smaller than $2e\cdot\ln n$ with probability greater than $1-\frac{2}{n}$.
Proof 5.5.
Let $u$ and $v$ be two nodes such that $u<v$, and let $X_{\text{up}}$ be the random variable equal to the increasing routing path length between $u$ and $v$ . Let $w\in[u,v]$, we note $1_{w}$ the indicator random variable equal to $1$ if $w$ is part of the increasing path between $u$ and $v$ (see proof of Theorem 5.2).
We first show that all $1_{w}$ are independent random variables. Let $w$ and $w^{\prime}$ be two nodes such that $u<w<w^{\prime}$, it holds:
$P[1_{w^{\prime}}=1\mid 1_{w}=1]$ $=P[h_{x}<h_{w^{\prime}}\forall x\in[w+1,w^{\prime}-1]\cap\max_{y\in[u,w]}h_{y}<h_{w}^{\prime}]$
$=P[h_{y}<h_{w}^{\prime}\forall y\in[u,w^{\prime}-1]]$
$=P[1_{w^{\prime}}=1]$
Now that we showed independence, we use the following Chernoff’s bound: For $Y$ the sum of independent Bernoulli random variables (a Poisson trial) and $\mu=\mathbb{E}[Y]$, $\forall\delta>0$ it holds that:
$P(Y\geq(1+\delta)\mu)<\left(\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\right)^{\mu}$
Replacing $Y=X_{\text{up}}=\sum_{w\in[u+1,v]}1_{w}$, $\mu=\ln n$ and $\delta=e-1$ gives:
$P(X_{\text{up}}\geq e\cdot\ln n)$ $<\frac{e^{\delta\cdot\ln n}}{(1+\delta)^{(1+\delta)\cdot\ln n}}$
$=\frac{n^{\delta}}{n^{(1+\delta)\cdot\ln(1+\delta)}}=\frac{1}{n}$
By symmetry, the same result holds for $X_{\text{down}}$. Finally, we obtain the desired result with a union bound:
$P(X\geq 2e\cdot\ln n)$ $\leq P(X_{\text{up}}\geq e\cdot\ln n\cup X_{\text{down}}\geq e\cdot\ln n)$
$\leq P(X_{\text{up}}\geq e\cdot\ln n)+P(X_{\text{down}}\geq e\cdot\ln n)$
$<\frac{2}{n}$
Theorem 5.6 (Expected Degree).
The expected degree of any node is lower than $4$.
Proof 5.7.
If $u$ and $w$ are two nodes, we define $N_{u}$ the random variable equal to the number of neighbors of $u$, and $1_{u,w}$ the indicator random variable equal to one if $u$ and $w$ are neighbors. It holds $N_{u}=\sum_{w\neq u}1_{u,w}$. Two nodes $u$ and $w$ with $u<w$ are neighbors if and only if $u$ and $w$ are the first and second highest nodes in $[u,w]$, hence it holds that $\mathbb{E}[1_{u,w}]=\frac{2}{|w-u|\cdot(|w-u|+1)}$. It therefore holds:
$\mathbb{E}[N_{u}]$ $=\sum_{w\in[1,n]\setminus\{u\}}\frac{2}{|w-u|\cdot(|w-u|+1)}$
$\leq 4\cdot\sum_{i=1}^{n/2}\frac{1}{i\cdot(i+1)}=4\cdot\sum_{i=1}^{n/2}\frac{1}{i}-\frac{1}{i+1}$
$=4\cdot(1-\frac{1}{n/2+1})\leq 4$
Theorem 5.8 (Maximum Degree with High Probability).
The maximum degree is lower than $4\ln n$ with probability greater than $1-\frac{1}{n^{3}}$.
Proof 5.9.
Let $u$, $w$ and $w^{\prime}$ be three different nodes. We name $1_{u,w}$ and $1_{u,w^{\prime}}$ the indicator random variables equal to $1$ if $u$ and $w$ are neighbors, and if $u$ and $w^{\prime}$ are neighbors, respectively. We first show that $1_{u,w}$ and $1_{u,w^{\prime}}$ are independent random variables: the proof is almost identical to the one of Theorem 5.4.
We first establish that each node taken individually has a degree lower than $\ln n$ with probability greater than $1-\frac{1}{n^{4}}$. Let $N_{u}$ be the random variable equal to the number of neighbors of $u$, we use the Chernov bound for Poisson trials used in the proof of Theorem 5.4:
$P[N_{u}\geq 4\ln n]$ $<\left(\frac{e^{(\ln n)-1}}{(\ln n)^{\ln n}}\right)^{4}=\frac{(n/e)^{4}}{n^{4\ln\ln n}}<\frac{1}{n^{4}}$
This last equality holds when $n$ is greater than a given constant, which is not a problem as we are interested in cases when $n$ is large. We conclude with a union bound: the probability that at least one node has a degree greater than $4\ln n$ is upper bounded by $\sum_{v\in V}P[N_{v}\geq 4\ln n]\leq\frac{1}{n^{3}}$ and the claim holds.
6 Learning-Augmented Continuous SLN
In this section, we present LASLiN, a learning-augmented Continuous SLN protocol where each node can obtain a prediction of the future demand matrix at any time.
We first show that, for general demands, any randomized, demand-oblivious P2P protocol can be off by a factor of $O(\log n/\log\log n)$ compared to an optimum static SLN. This result provides a comparison point around which the consistency and robustness metrics of demand-aware algorithms can be evaluated.
Theorem 6.1.
We consider any randomized P2P protocol (Definition 5.1) that is ideal (Definition 2.4) and not learning-augmented (Definition 2.7). Then, there exists a demand matrix $W$ such that the considered protocol has a cost of $\Omega(\log n/\log\log n)\cdot|W|$ in expectation while an optimal static SLN protocol pays $|W|$, where $|W|$ is the total sum of the elements of $W$.
Proof 6.2.
As the considered protocol is ideal, there exists a constant $c$ such that the generated network has a maximum degree lower than $O(\ln^{c}n)$ with probability greater than 1O(1/n)1-O(1/n). Let $u$ be a node. We first cover the case when the maximum degree is lower than $\ln^{c}n$, all the considered probabilities and expected values are therefore conditional. In this case, we consider the disk centered around $u$ with smallest radius such that it contains no less than half of the nodes. The maximum distance between $u$ and a node in that disk is no lower than $O(\log n/\log\log n)$ as the network has a polylogarithmic maximum degree — this result can be easily obtained by upper bounding all the degrees by $O(\log^{c}n)$. There exists another node $v$ such that $v$ is at the border or outside of the disc with probability at least $1/2$. We consider the demand matrix that is zero everywhere except at the coefficient with row $u$ and column $v$. An optimal SLN protocol can clearly pay a cost of $1$, the considered randomized SLN therefore has an approximation ratio worse of $\omega(\log n/\log\log n)$ for that demand matrix.
Finally, we cover the cases where the maximum degree is higher than $\ln^{c}n$: That case only contributes with an additive constant to the distance between $u$ and $v$ as a distance is lower than $n$. We considered all the cases and the claim follows.
Protocol 2 (LASLiN).
LASLiN is a randomized, continuous SLN that sets its heights by adding the input, integer prediction in $1,\lceil\ln n\rceil]$, and a real number in $(0,1)$ chosen uniformly at random.
Theorem 6.3.
LASLiN is a randomized, Continuous SLN that is $O(1)$-consistent and $O(\log^{2}n)$-robust.
Proof 6.4.
We start by proving that the maximum degree in LASLiN is lower than $4\ln^{2}n$ with high probability. Let $u$ be a node, let $p_{u}\in[1,\lceil\ln n\rceil]$ be the predicted height of $u$ and let $r_{u}\in(0,1)$ be the uniform random noise added on top of $p_{u}$. We partition the neighbors of $u$ into $p_{u}$ subsets called $S(1),S(2)\dots S(p_{u}+1)$: For all $i\in[1,p_{u}]$, $S(i)$ contains the neighbors with height between $i$ and $i+1$, the last set $S(p_{u}+1)$ contains the neighbors with height greater than $p_{u}+1$. According to Theorem 5.8, each set from $S(1)$ to $S(p_{u})$ contains fewer than $4\ln n$ nodes with probability greater than $1-\frac{1}{n^{4}}$ (this is the result before the last paragraph in the proof) — the set $S(p_{u}+1)$ contains at most two nodes. We conclude with a union bound that $u$ has fewer than $4\ln^{2}n$ with probability greater than $1-\frac{1}{n^{3}}$, and that the maximum degree is lower than $4\ln^{2}n$ with probability greater than $1-\frac{1}{n^{2}}$ also with a union bound.
The $O(1)$-consistency is direct as the comparative optimal static SLN has its heights within a $O(\log n)$ factor without loss of generality as its maximum degree is bounded by $O(\log n)$ by assumption. We now prove the robustness claim. A routing path between any two nodes $u$ and $v$ has two phases: first ascending (the heights always increase) and then descending (the heights always decrease). For both phase, the number of different encountered integer heights is upper-bounded by $O(\log n)$. For each encountered integer height, the expected routing path length is upper-bounded by $O(\log n)$ (Theorem 5.2). We finish the proof with an upper bound of each distance in the total weighted path lengths: $\sum_{u,v\in V}W(u,v)\cdot d(u,v)\leq O(\log^{2}n)\cdot|W|$ and the optimal static SLN pays at least $|W|$.
We now define the notion of super-sparse matrix, which indicates a concentration of the weights of the matrix on a very small (logarithmic) number of coefficients.
Definition 6.5 (super-sparse matrix).
A matrix $W\in M_{n}(\mathbb{R})$ is super-sparse if there exists a constant $c\geq 1$ and a restricted subset of node pairs $S\subseteq V^{2}$ such that:
$|S|$ $\leq O(\log^{c}n)\qquad\text{ and }$
$\sum_{(u,v)\in S}W(u,v)$ $\geq\left(1-O(1/\log n)\right)\cdot|W|$
Theorem 6.6.
Assume that the demand matrix $W$ is super-sparse. We consider the protocol Binary LASLiN (B-LASLiN) where the input prediction $p$ can only have two values, $1$ or $2$. B-LASLiN is $O(\log\log n)$-consistent and $O(\log n)$-robust.
In asymptotics, the performances are therefore always better than any demand-oblivious protocol, with an exponential improvement in case of perfect predictions.
Proof 6.7.
We start by proving that the maximum degree of each node is lower than $O(\log n)$ with high probability. As seen in the proof of Theorem 6.3, each node $u$ has fewer than $p_{u}\cdot\ln n$ neighbors with high probability, the claim follows since $p_{u}\leq 2$.
We prove that the expected total weighted path length is lower than $O(\log n)\cdot|W|$ regardless of the input predictions. The proof is similar to the proof of Theorem 6.3 except, as the predictions are upper-bounded by a constant, we avoid the additional $\log n$ factor.
Finally, we prove that the expected total weighted path length is lower than $O(\log\log n)\cdot|W|$. We consider the following set of predictions: let $S$ be the node pairs from Definition 6.5, we give prediction $2$ to the nodes that have at least one coefficient from $S$ in their row or column in $W$, otherwise $1$. We call $V_{2}$ ($V_{1}$) the set of nodes that have prediction 22 (11). As $W$ is super sparse, it holds that $|V_{2}|\leq O(\log^{c}n)$.
$\sum_{u,v}W(u,v)\cdot d(u,v)$ $=\sum_{(u,v)\in S}W(u,v)\cdot d(u,v)+\sum_{(u,v)\in V^{2}\setminus S}W(u,v)\cdot d(u,v)$
$\leq\sum_{u,v\in V_{2}}W(u,v)\cdot d(u,v)+\sum_{(u,v)\in V^{2}\setminus S}W(u,v)\cdot d(u,v)$
$\leq O(\log\log n)\cdot|W|+O\left(\frac{1}{\log n}\right)\cdot|W|\cdot\ln n$
$=O(\log\log n)\cdot|W|$
To go from line 3 to 4 we used (1) Theorem 5.2 on the routing paths between the nodes of $V_{2}$ — there are $O(polylog(n))$ of those nodes, they have the largest heights of the SLN, so a path has expected length $\ln O(polylog(n))=O(\log\log n)$ — and on the routing paths from nodes in $V^{2}\setminus S$, and (2) we used Definition 6.5 to upper-bound the sum of coefficients in $W$ outside of $S$.
Figure 6: Comparison of the costs of multiple P2P networks against different demand matrix distributions. The demand matrices are generated based on the Zipfian distribution with varying skewness parameter $\zeta\in\{1.1,1.2\dots 2.0\}$, each coefficient is drawn independently and identically distributed. For each demand matrix distribution, we compare the cost of three P2P networks: the optimal static SLN, LASLiN, and the demand-oblivious Uniform SLN.
7 Experimental Evaluation
Demand matrices
In our experiments, we generate communication demand based on the Zipfian distribution [1], which receives a parameter $\zeta>0$, that indicates the distribution’s skewness. Roughly speaking, lower values of $\zeta$ result in a sparser demand matrix. The number of nodes in all experiments was set to $n=32$, and each experiment was repeated $30$ times.
Noise generation
As common in the literature [20], we study the robustness of our algorithms by introducing noise in the predicted data. The most used type of noise is additive, often gaussian [26] or lognormal [20]. This type of noise however almost did not impact the performance of our algorithms. Instead, we test two other models: the demand error recomputes a portion (x-axis) of the coefficients with the same Zipfian distribution, and the unit heights enforce that each node can receive a predicted height of $1$ with a certain probability (x-axis).
Results
In Figure 6 we evaluate the consistency of LASLiN in scenarios where the demand follows the Zipfian distribution, with varying parameter $\zeta\in[1.1,2]$, and where the prediction advice is perfect (i.e., there is zero noise in the demand matrix). We compare the total communication cost (as defined in (2.5)) of LASLiN to that of the static optimum SLN and the demand-oblivious SLN (Uniform) baselines. We can see that, in contrast to the demand-oblivious SLN, LASLiN obtains close-to-optimum total communication cost in all workloads, regardless of their spatial locality, i.e., LASLiN is consistent.
In Figure 7 we evaluate the robustness of LASLiN in scenarios with increasing demand noise (x-axis). We compare the total cost of LASLiN to that of the static optimum SLN, the demand-oblivious (Uniform) SLN, and the SLN baselines. We can observe that the SLN performs increasingly worse (has higher cost) the greater the amount of change (noise) in the input demand matrix. LASLiN, on the other hand, achieves consistently lower communication cost (than the former), regardless of the type of noise and regardless of the prediction fault probability, i.e., LASLiN is robust relative to both type and magnitude of prediction errors.
Figure 7: Cost of different P2P networks under varying noise amplitudes on the demand matrix predictions. Simulation results: robustness ($\zeta=1.2$)
8 Related Work
Improved designs of P2P networks. P2P networks have been the focus of computer scientists for a long while (see [19] for a detailed survey). In terms of P2P networks with constant degree, the notable contribution of [23] that extends the work of [13], giving a general P2P construction such that for a network with degree $d$ and $n$ nodes, you can have guaranteed path length of $\Theta(\log_{d}n)$. Lately, improving the design of networks given prior information about communication demand has grabbed the attention of researchers. In particular, the work of [7] set the scene for demand-aware network design, which was then explored more with particular focus on datacenter networking [11, 24, 12].
Works closest to ours are by [6, 10] in which the authors discuss a self-adjusting variant of skip list network: one that adjusts over time based on the incoming demand, which is in contrast with our work that focuses on an optimal skip list network with static height given the communication demand as the input. This is in particular important, as self-adjustments can introduce additional cost (e.g., in terms of delay) to the system. Also, the work of [21] discussed how a skip-list construction can be used for peer-to-peer communications, however they do not provide any theoretical upper bound for the communication time. Lastly, we point out to other extensions of skip list, namely Skip Graph [3] and SkipNet [16] that, to facilitate P2P communications, consider expected degree to be $O(\log n)$, in contrast to our design, which has a constant degree in expectation.
Learning-augmented algorithms. Given the rise of machine learning algorithms (that are prone to inherent unbounded failures), and after the pioneering work of [26] on ski rental & job scheduling, the focus has been shifted on studying consistency and robustness of algorithms augmented with the learned advice. Such an approach has been used in various applications so far222see a full list of recent works in the following website: algorithms-with-predictions.github.io, for example caching [20], online TSP [15], list update [8].
Some of the more recent works focus on improving single-source search data structures. Initially, the works of [18, 9] provided learning-augmented search trees. The works that are closest to our work are by [14, 30], which studied learning-augmented single-source skip lists. In contrast, we consider a more challenging model, where the focus is on pairwise communications rather than communication from a single source. Recently, we became aware of a not peer-reviewed manuscript [2] that aims to augment P2P networks with learned advice; however, this paper looks into improving self-stabilization and recovery of P2P networks, not improving communication delay.
9 Conclusion
In this work, introduced a continuous variant of Skip List Network (SLN), Continuous SLN. This variant is arguably easier to manipulate and analyze. In addition we show that we can use it to reach strong robustness properties even in an adversarial, fully distributed setting. Lastly, through experimental evaluation, we observed that our suggested algorithms outperform the theoretical results. As a future work, one can think of an alternative algorithm for OSSLN with a lower computational complexity.